# Enhanced Electromagnetic Memory

### v1

@@ Acknowledgements.
I am very grateful to Vitor Cardoso and David Garfinkle for valuable input and comments on a first version of the manuscript. JZ is supported by funding from the Swiss National Science Foundation grant 222346. The Center of Gravity is a Center of Excellence funded by the Danish National Research Foundation under grant No. 184.
@@ Appendix A Full Expressions of Electromagnetic Memory
In this appendix, we explicitly write out the solutions for the electromagnetic memory in Eqs. (15) and (16) in terms of polarization content of the vector potential
$A^{T}_{\hat{\theta}}\equiv A_{i}\hat{\theta}_{i}\,,\quad A^{T}_{\hat{\phi}}\equiv A_{i}\hat{\phi}_{i}\,,$ (25)
Here, we have defined the standard spacial transverse basis vectors
@@
[…]
that is completed through the longitudinal direction
$\hat{\mathbf{n}}(\Omega)=(\sin\theta\cos\phi,\,\sin\theta\sin\phi,\,\cos\theta)\,,$ (27)
such that $\delta_{ij}=n_{i}n_{j}+u_{i}u_{j}+v_{i}v_{j}$.
@@
$\tilde{A}^{T}_{\hat{\theta}}(\psi)$ $=\cos\psi A^{T}_{\hat{\theta}}-\sin\psi\tilde{A}^{T}_{\hat{\phi}}\,,$ (28)
$\tilde{A}^{T}_{\hat{\phi}}(\psi)$ $=\sin\psi A^{T}_{\hat{\theta}}+\cos\psi\tilde{A}^{T}_{\hat{\phi}}\,,$ (29)
and the associated polarization angle choice that fulfills the conventions above is
@@
${}_{\scriptscriptstyle{\text{(I)}}}\tilde{A}^{T}_{\hat{\theta}}(\psi_{\hat{\theta}})$ $=\frac{q\,\mu v_{q}}{4\pi r}\Theta(u-u_{q})\Gamma(\Omega,\Omega_{q})\,,$ (31a)
${}_{\scriptscriptstyle{\text{(I)}}}\tilde{A}^{T}_{\hat{\phi}}(\psi_{\hat{\theta}})$ $=0\,,$ (31b)
${}_{\scriptscriptstyle{\text{(II)}}}\tilde{A}^{T}_{\hat{\theta}}(\psi_{\hat{\theta}})$ $=\frac{q\,\mu v_{q}}{4\pi r}[1-\Theta(u-u_{q})]\Gamma(\Omega,\Omega_{q})\,,$ (31c)
${}_{\scriptscriptstyle{\text{(II)}}}\tilde{A}^{T}_{\hat{\phi}}(\psi_{\hat{\theta}})$ $=0\,,$ (31d)
@@
The total amplitude of the electromagnetic memory in each of the cases C $=$ (I), (II) is then defined as
@@
In this case, the total memory electromagnetic amplitude within the vector potential [Eq. (33)] reads
$\Delta A^{T}(r,\Omega,\{0,0\})$ =$=\begin{cases}\frac{q\mu v_{q}}{4\pi r}\frac{\sin\theta}{1-\frac{v_{q}}{v_{\text{em}}}\cos\theta},&\text{if $\frac{v_{q}}{v_{\text{em}}}\cos\theta<1$}.\\ \frac{q\mu v_{q}}{4\pi r}\frac{-\sin\theta}{1-\frac{v_{q}}{v_{\text{em}}}\cos\theta},&\text{if $\frac{v_{q}}{v_{\text{em}}}\cos\theta>1$}.\end{cases}$ (35)
@@ Appendix B The Energy Spectrum of Electromagnetic Memory
$E^{i}_{T}(u,r,\Omega)$ =$=\begin{cases}\frac{-q\mu v_{q}}{4\pi r}\frac{D(u-u_{q})\,\perp^{ij}\hat{n}_{q}^{j}}{1-\frac{v_{q}}{v_{\text{em}}}\hat{\mathbf{n}}\cdot\hat{\mathbf{n}}_{q}},&\text{if $\hat{\mathbf{n}}\cdot\hat{\mathbf{n}}_{q}<\frac{v_{\text{em}}}{v_{q}}$},\\ \frac{q\mu v_{q}}{4\pi r}\frac{D(u-u_{q})\,\perp^{ij}\hat{n}_{q}^{j}}{1-\frac{v_{q}}{v_{\text{em}}}\hat{\mathbf{n}}\cdot\hat{\mathbf{n}}_{q}},&\text{if $\hat{\mathbf{n}}\cdot\hat{\mathbf{n}}_{q}>\frac{v_{\text{em}}}{v_{q}}$},\end{cases}$ (37)
@@
$E_{\text{tot}}\simeq\frac{\mu q^{2}}{8\pi^{3/2}}\frac{\gamma_{q}m_{q}c^{2}}{\hbar}v_{\text{em}}^{2}\left(\frac{v_{\text{em}}}{v_{q}}\ln\left[\frac{1+v_{q}/v_{\text{em}}}{1-v_{q}/v_{\text{em}}}\right]-2\right)\,.$ (42)
In this estimate of the energy emitted in electromagnetic memory radiation, we have used that for the emission of a single charged particle of mass $m_{q}$, the emission time scale $\Delta u$ can be estimated through the Heisenberg uncertainty principle to be of order
$\Delta u\sim\frac{\hbar}{E_{q}}=\frac{\hbar}{\gamma_{q}m_{q}c^{2}}\,.$ (43)
As a conclusion, the magnification in the memory amplitude, directly translates into an enhancement of the total emitted energy, with the biggest support coming from the low frequency spectrum and the total radiative energy diverges as $v_{q}\rightarrow v_{\text{em}}$.
@@ Appendix C Emission Along Lightcone Directions
$z_{c}$ =±$=\pm\sqrt{t^{2}v_{q}^{2}-\left\{tv_{\text{em}}+\frac{t^{2}(v_{q}-v_{\text{em}})(v_{q}+v_{\text{em}})}{2rv_{\text{em}}/c}\right\}^{2}}\,.$ (47b)
As expected, there is only a real solution if $v_{q}\geq v_{\text{em}}$. This can be seen more clearly by explicitly computing the normalized spacial projection of the emission direction of the charge $\hat{\mathbf{n}}^{c}_{q}$ that corresponds to an emission along the lightcone directions $\partial_{t}x_{c}|_{t=0}$ and $\partial_{t}z_{c}|_{t=0}$
$\hat{\mathbf{n}}_{q}^{c}=(v_{\text{em}}/v_{q},0,\sqrt{1-v_{\text{em}}^{2}/v_{q}^{2}})\,.$ (48)
The critical direction of charge emission $\hat{\mathbf{n}}^{c}_{q}$ along the intersection of the past causal light cone defined through a direction of memory evaluation $\hat{\mathbf{n}}$ in Eq. (46) is therefore characterized by a scalar product satisfying
$\hat{\mathbf{n}}\cdot\hat{\mathbf{n}}^{c}_{q}=\frac{v_{\text{em}}}{v_{q}}\,.$ (49)
This condition precisely coincides with the critical directions of diverging memory amplitude in Eq. (17) identified in section II.3. The corresponding geometrical configuration is depicted in Fig. 3.
Figure 3: Causal cone (blue) of emission of a charge $q$ at a speed $1>v_{q}>v_{\text{em}}$ with $c=1$. Shown are only two spacial directions $x$ and $z$ and the time direction $t$. The cone intersects a past null cone of a memory event (orange) admitting critical directions of emission that give rise to a formally unbound electromagnetic memory buildup. Shown are also the spacial projections of the cone (dashed blue) and the intersection line (dashed red), as well as representatives of the directions $\hat{\mathbf{n}}$ of memory evaluation and $\hat{\mathbf{n}}_{q}$ of charge emission. The angle $\hat{\mathbf{n}}\cdot\hat{\mathbf{n}}^{c}_{q}=\cos\alpha$ between the two directions precisely corresponds to the critical direction identified in Eq. (17) $\alpha=\arccos[v_{\text{em}}/v_{q}]$.

### v2

@@ Acknowledgements.
I am very grateful to Vitor Cardoso and David Garfinkle for valuable input and comments on a first version of the manuscript. I also want to thank an anonymous referee for constructive comments and in particular for pointing out the potential application of the work on the tail effects associated to the sub-leading soft theorem. JZ is supported by funding from the Swiss National Science Foundation (Grant No. 222346). The Center of Gravity is a Center of Excellence funded by the Danish National Research Foundation under Grant No. 184.
@@ Appendix A Full Expressions of Electromagnetic Memory
In this appendix, we explicitly write out the solutions for the electromagnetic memory in Eqs. (15) and (16) in terms of polarization content of the vector potential $A^{T}_{\hat{\theta}}\equiv A_{i}\hat{\theta}_{i}$ and $A^{T}_{\hat{\phi}}\equiv A_{i}\hat{\phi}_{i}$. The standard spacial transverse and associated longitudinal basis vectors read
@@
$\hat{\bm{\phi}}(\Omega)$ $=(-\sin\phi,\,\cos\phi,\,0)\,$, (25b)
$\hat{\mathbf{n}}(\Omega)$ $=(\sin\theta\cos\phi,\,\sin\theta\sin\phi,\,\cos\theta)\,$, (25c)
satisfying $\delta_{ij}=n_{i}n_{j}+u_{i}u_{j}+v_{i}v_{j}$.
@@
$\tilde{A}^{T}_{\hat{\theta}}(\psi)$ $=\cos\psi A^{T}_{\hat{\theta}}-\sin\psi\tilde{A}^{T}_{\hat{\phi}}\,$, (26a)
$\tilde{A}^{T}_{\hat{\phi}}(\psi)$ $=\sin\psi A^{T}_{\hat{\theta}}+\cos\psi\tilde{A}^{T}_{\hat{\phi}}\,.$ (26b)
The angle $\psi$ that fulfills the conventions above is
@@
[…]
@@
The total amplitude of the electromagnetic memory in each of the cases C == (I), (II) is then defined as
@@
In this case, the total electromagnetic memory amplitude within the vector potential [Eq. (30)] reads
$\Delta A^{T}(r,\Omega,\{0,0\})$ $=\begin{cases}\frac{q\mu v_{q}}{4\pi r}\frac{\sin\theta}{1-\frac{v_{q}}{v_{\text{em}}}\cos\theta},&\text{if $\frac{v_{q}}{v_{\text{em}}}\cos\theta<1$}.\\ \frac{q\mu v_{q}}{4\pi r}\frac{-\sin\theta}{1-\frac{v_{q}}{v_{\text{em}}}\cos\theta},&\text{if $\frac{v_{q}}{v_{\text{em}}}\cos\theta>1$}.\end{cases}$ (32)
@@ Appendix B The Energy Spectrum of Electromagnetic Memory
$E^{i}_{T}(u,r,\Omega)$ $=\begin{cases}\frac{-q\mu v_{q}}{4\pi r}\frac{D(u-u_{q})\,\perp^{ij}\hat{n}_{q}^{j}}{1-\frac{v_{q}}{v_{\text{em}}}\hat{\mathbf{n}}\cdot\hat{\mathbf{n}}_{q}},&\text{if $\hat{\mathbf{n}}\cdot\hat{\mathbf{n}}_{q}<\frac{v_{\text{em}}}{v_{q}}$},\\ \frac{q\mu v_{q}}{4\pi r}\frac{D(u-u_{q})\,\perp^{ij}\hat{n}_{q}^{j}}{1-\frac{v_{q}}{v_{\text{em}}}\hat{\mathbf{n}}\cdot\hat{\mathbf{n}}_{q}},&\text{if $\hat{\mathbf{n}}\cdot\hat{\mathbf{n}}_{q}>\frac{v_{\text{em}}}{v_{q}}$},\end{cases}$ (34)
@@
$E_{\text{tot}}\simeq\frac{\mu q^{2}}{8\pi^{3/2}}\frac{\gamma_{q}m_{q}c^{2}}{\hbar}v_{\text{em}}\left(\frac{v_{\text{em}}}{v_{q}}\ln\left[\frac{1+v_{q}/v_{\text{em}}}{1-v_{q}/v_{\text{em}}}\right]-2\right)\,.$ (39)
In this estimate of the energy emitted in electromagnetic memory radiation, we have used that for the emission of a single charged particle of mass $m_{q}$, the emission time scale $\Delta u$ can be estimated through the Heisenberg uncertainty principle to be of order $\Delta u\sim\frac{\hbar}{E_{q}}=\frac{\hbar}{\gamma_{q}m_{q}c^{2}}$. As a conclusion, the magnification in the memory amplitude, directly translates into an enhancement of the total emitted energy, with the biggest support coming from the low frequency spectrum and the total radiative energy diverges as $v_{q}\rightarrow v_{\text{em}}$.
@@ Appendix C Emission Along Lightcone Directions
$z_{c}$ $=\pm\sqrt{t^{2}v_{q}^{2}-\left\{tv_{\text{em}}+\frac{t^{2}(v_{q}-v_{\text{em}})(v_{q}+v_{\text{em}})}{2rv_{\text{em}}/c}\right\}^{2}}\,.$ (43b)
Figure 3: Causal cone of emission of a charge $q$ at a speed $1>v_{q}>v_{\text{em}}$ with $c=1$ (blue) intersecting a past null cone of a memory event (orange) admitting critical directions of memory enhancement. The angle $\alpha$ precisely corresponds to the critical angle identified in Eq. (17).
As expected, there is only a real solution if $v_{q}\geq v_{\text{em}}$. This can be seen more clearly by explicitly computing the normalized spacial projection of the emission direction of the charge $\hat{\mathbf{n}}_{q}^{c}=(v_{\text{em}}/v_{q},0,\sqrt{1-v_{\text{em}}^{2}/v_{q}^{2}})$ that corresponds to an emission along the lightcone directions $\partial_{t}x_{c}|_{t=0}$ and $\partial_{t}z_{c}|_{t=0}$. The critical direction of charge emission $\hat{\mathbf{n}}^{c}_{q}$ along the intersection of the past causal light cone defined through a direction of memory evaluation $\hat{\mathbf{n}}$ in Eq. (42) is therefore characterized by a scalar product satisfying $\hat{\mathbf{n}}\cdot\hat{\mathbf{n}}^{c}_{q}=\frac{v_{\text{em}}}{v_{q}}$. This condition precisely coincides with the critical directions of diverging memory amplitude in Eq. (17) identified in section II.3. The corresponding geometrical configuration is depicted in Fig. 3.